OpenStudy (lifeisadangerousgame):

How do I add 6/7 x 3?

OpenStudy (lifeisadangerousgame):

multiply I mean

OpenStudy (ragingsquirrel):

Turn 6/7 into a number.

OpenStudy (lifeisadangerousgame):

sorry, I kinda messed up the equation, that's not what I meant

hero (hero):

You shouldn't have....

OpenStudy (ragingsquirrel):

P_P Way to be harsh on a noob...

OpenStudy (lifeisadangerousgame):

lol sorry, That's what the model does, but I want to know how to do that on my problemn

OpenStudy (lifeisadangerousgame):

What I want is to turn this: y + 1 = 6/7(x + 3) into one of these: A. –6x + 5y = 13 B. –6x +7y =–11 C. –6x + 7y = 11 D. –7x + 6y =–11

hero (hero):

y + 1 = 6/7(x + 3) y + 1 = 6x/7 + 18/7 y = 6x/7 + 18/7 - 1 y = 6x/7 + 18/7 - 7/7 y = 6x/7 + 11/7 y = (6x + 11)/7 7y = 6x + 11 7y = 11 + 6x -11 = -7y + 6x -6x + 7y = 11

OpenStudy (ragingsquirrel):

Number 3.

OpenStudy (lifeisadangerousgame):

Thanks you!